# Theory Uncertainties in the Extraction of $\alpha_{s}$ from Drell-Yan at Small Transverse Momentum

### v1

@@ A.2 N4LL approximation uncertainties
Different nomenclatures for resummation orders involving Sudakov double-logarithms appear in the literature. A common nomenclature, which we also use here, is such that at NnLL one includes the $n$-loop evolution (i.e. $n$-loop noncusp and $(n+1)$-loop cusp anomalous dimensions) together with $(n-1)$-loop boundary conditions, which in our notation are the fixed-order hard, beam, and soft functions. At NnLL order the boundary conditions are included instead at $n$-loop order, i.e., one order higher than at NnLL.
The NnLL nomenclature used in Refs. [33, 30] on the other hand includes $n$-loop boundary conditions so it corresponds more closely [^12] It might not be identical to NnLL as implemented in SCETlib due to various choices of how the different ingredients are included in detail, leading to differences that are considered to be of higher order in one or the other counting. to what we would call NnLL and to avoid confusion we will always use the latter notation. The analysis in Ref. [33] is performed at the so-called approximate N4LL (“N4LLa”) order obtained in Ref. [30], which in our nomenclature therefore translates to an approximate N4LL order. There are two types of approximation involved: First, the 4-loop evolution is approximate, because the required 4-loop splitting functions and 5-loop cusp anomalous dimensions are only known approximately, which is expected to cause more minor effects and is not discussed here further. [^13] Although the available approximations for 4-loop splitting functions are increasingly accurate for phenomenology [101], their effect could in principle be of some relevance and deserves further study. More importantly, another approximation comes from the fact that the 4-loop beam and soft functions required at this order are completely unknown.
@@
Toward higher orders in $q_{T}$ resummation, the fixed-order boundary conditions become the dominant effect while the anomalous dimensions become less important. For example, the SCETlib results for the $q_{T}$ spectrum at N3LL and N4LL, which only differ by the inclusion of higher anomalous dimensions, are almost indistinguishable in both central value and scale variation uncertainties [32]. [^14] This applies when using common PDFs. Switching from NNLO to approximate N3LO PDFs, as also formally required for the PDF evolution when switching from N3LL to N4LL, does make a noticeable difference. Furthermore, the fixed-order beam and soft functions are more important than the hard function, since the latter is genuinely evaluated at a high scale $\sim m_{Z}$, while the former two are genuinely evaluated at the low scale scale $\sim q_{T}\sim 1/b_{T}$. [^15] The presence of a perturbative series evaluated at the low scale $\mu\sim q_{T}$ is a universal feature of $q_{T}$ resummation and not specific to a particular resummation framework. In DYturbo the corresponding contributions are superficially evaluated at a high scale but the evolution factor includes additional contributions that transport them back to their natural low scale. The 4-loop hard function is known, while the 4-loop beam and soft functions are currently unknown. This means that the approximate N4LL order used in Refs. [30, 33] cannot be regarded as a close approximation of the true N4LL. Instead, it is a largely incomplete order, as the most important perturbative contributions for this order are missing.
In Refs. [30, 33] the additional uncertainty due to this approximation is estimated to be $0.1-0.2\%$ for the $q_{T}$ spectrum with a resulting negligible ±$\pm 0.04\times 10^{-3}$ effect on $\alpha_{s}(m_{Z})$. However, given that the most important perturbative ingredients at N4LL are missing, the associated uncertainty should be roughly of comparable size as the perturbative uncertainty of the previous N3LL order, which after all is meant to estimate the uncertainty due to exactly these (and other less important) N4LL contributions. The scale variations at N3LL may of course only provide a rough measure and a more explicit estimate of the individual missing terms might be more precise, but it is inconceivable that their estimated effect would be be negligibly small in comparison. Moreover, the estimate is 10 times smaller than the scale variations at the approximate N4LL order itself, which estimate the uncertainty due to even higher $5$-loop corrections, which is clearly inconsistent.
@@
The dominant effect of the beam and soft functions is also evident in our TNP studies, where as expected they cause by far the largest uncertainties. From our TNP scanning at N3+1LL, we know that the 3-loop beam and soft functions each cause an effect of order $\sim 1\times 10^{-3}$ on $\alpha_{s}(m_{Z})$. We might expect the corresponding uncertainties due to the 4-loop beam and soft functions to be between two and a few times smaller than that, but not 30 to 40 times smaller. We can obtain a more realistic estimate of the impact of the missing 4-loop beam and soft functions from their associated TNPs at N4+0LL (see appendix B.2). [^16] Ideally, we would use N4+1LL, which we have not yet available, but N4+0LL provides a sufficient approximation for our purposes here. From the analogous TNP scanning at this order we find a combined perturbative uncertainty on $\alpha_{s}(m_{Z})$ due to the missing 4-loop beam and soft functions of $0.75\times 10^{-3}$, as given in table 2, so indeed a reduction by roughly a factor of 2 compared to N3+1LL in line with expectation. It is also twice as large as the quoted scale-variation uncertainty at the highest approximate working order.
@@ A.3 Effects due to finite bottom and charm quark masses
Ref. [33] uses a massless 5-flavor description by default. The uncertainties due to neglecting quark-mass effects are estimated by changing to a variable-flavor massless description in various parts of the calculation and by varying the associated threshold scales. From the above discussion, this can only serve as a rough indication of the expected order of magnitude of mass effects but it may or may not reflect the actual size of the effects. Another estimate comes from including the effect of final-state gluon splitting into massive quark pairs. Unfortunately, Ref. [33] provides no further information than that, so it is unclear to us to what extent this provides a complete description of secondary mass effects. The final uncertainty on $\alpha_{s}(m_{Z})$ due to mass effects is taken as the envelope of all observed differences yielding ×${}^{+0.40}_{-0.29}\times 10^{-3}$.
@@ A.4 Nonperturbative model
This model originates from Ref. [89], and while it has traditionally been used in the past, it has several limitations that are potentially important at the high level of precision required here. First, the model does not account for any flavor or $x$ dependence in the nonperturbative TMD parameters. As discussed in section 3.3.2, since the fit to real data includes multiple rapidity bins, the nonperturbative TMD model should at minimum allow for an effective rapidity dependence.
@@ B.1 Nonsingular power corrections
We perform a set of Asimov tests following the procedure in section 4.2.3. For simplicity, we do not fit the nonperturbative parameters here, since doing so does not change our conclusions. The nominal perturbative uncertainty for this case obtained in section 4.2.3 is $\Delta_{\rm pert}=_{-0.50}^{+0.47}\times 10^{-3}$. The Asimov data are now defined as the N4LL$+$NLO1 prediction at $\alpha_{s}(m_{Z})=0.118$, which we fit with three different theory models that differ in the nonsingular terms they include. The results are shown in table 3. In the case where no nonsingular term is included (top row), we find an uncertainty $\Delta_{\rm pert}=0.41\times 10^{-3}$, slightly smaller than the nominal one. Including the LO1 and NLO1 terms (middle and bottom rows), the uncertainty increases again marginally and closely reproduces the nominal one. We conclude that under the assumptions and setup of our pseudodata analysis, the nonsingular contributions do not affect the uncertainty of $\alpha_{s}(m_{Z})$. They can therefore be safely neglected for our purposes. However, as also seen in table 3, neglecting the nonsingular corrections in the theory model induces a substantial bias in the central value of $\alpha_{s}(m_{Z})$, which will necessitate their inclusion in the fit to real data.

### v2

@@ A.2 Different N4LL countings and approximation uncertainties
Different nomenclatures for resummation orders involving Sudakov double-logarithms appear in the literature. A common nomenclature, which we also use here, is such that at NnLL one includes the $n$-loop evolution (i.e. $n$-loop noncusp and $(n+1)$-loop cusp anomalous dimensions) together with (n1)(n-1)-loop fixed-order boundary conditions of the hard, beam, and soft functions in our notation. At NnLL order the boundary conditions are included at one order higher, i.e., at $n$-loop order and as they appear at Nn+1LL. As a result, the perturbative precision at NnLL is often close to Nn+1LL and thus much higher than NnLL. This is summarized in table 3.
The NnLL nomenclature used in Refs. [33, 30] also includes $n$-loop boundary terms, and for clarity we will refer to it as NnLL here. According to Ref. [30], this N4LL supposedly provides a reliable approximation of N4LL, with the implied higher perturbative precision compared to standard N4LL. However, there is a crucial difference in how the boundary terms are included at NnLL, making it rather different from N4LL and with a perturbative precision more comparable to standard N4LL, as indicated in table 3 and as we will discuss now.
boundary conditions anomalous dimensions
resummation order $h$ $\tilde{b}$, $\tilde{s}$ $\gamma^{\mu},\gamma^{\nu},P_{ij}$ $\Gamma_{\mathrm{cusp}},\beta$
N3LL $\alpha_{s}^{3}(Q)$ $\alpha_{s}^{3}(b_{0}/b_{T})$ $\alpha_{s}^{3}$ $\alpha_{s}^{4}$
N4LL $\alpha_{s}^{3}(Q)$ $\alpha_{s}^{3}(b_{0}/b_{T})$ $\alpha_{s}^{4}$ $\alpha_{s}^{5}$
N4LL $\alpha_{s}^{4}(Q)$ $\alpha_{s}^{3}(b_{0}/b_{T})+\alpha_{s}^{4}(Q)$ $\alpha_{s}^{4}$ $\alpha_{s}^{5}$
N4LL $\alpha_{s}^{4}(Q)$ $\alpha_{s}^{4}(b_{0}/b_{T})$ $\alpha_{s}^{4}$ $\alpha_{s}^{5}$
N5LL $\alpha_{s}^{4}(Q)$ $\alpha_{s}^{4}(b_{0}/b_{T})$ $\alpha_{s}^{5}$ $\alpha_{s}^{6}$
Table 3: Order counting conventions. Here, N4LL refers to what is called N4LL in Refs. [33, 30]. In practice, the first three rows have a comparable perturbative precision, as have the last two rows. See the text for further details.
Let us first consider the soft function. As discussed in section 3.2, the full soft function $\tilde{S}(b_{T},\mu,\nu)$ is a perturbative series in $\alpha_{s}(\mu)$ and contains Sudakov logarithms $\ln^{n}(\mu b_{T}/b_{0})$ and $\ln^{n}(\nu b_{T}/b_{0})$. Its resummation proceeds by evaluating it at its canonical scales $\mu_{S}^{\rm can}=\nu_{S}^{\rm can}=b_{0}/b_{T}$ [see Eqs. (3.15) and (3.8)], which defines its boundary condition,
$\tilde{s}[\alpha_{s}(b_{0}/b_{T})]\equiv\tilde{S}(b_{T},\mu=b_{0}/b_{T},\nu=b_{0}/b_{T})=\sum_{n=0}\tilde{s}_{n}\,\biggl[\frac{\alpha_{s}(b_{0}/b_{T})}{4\pi}\biggr]^{n}\,$, (A.1)
where the $\tilde{s}_{n}$ are the so-defined $n$-loop soft boundary coefficients. The boundary series in Eq. (A.1) is by construction free of large logarithms and can thus be evaluated at fixed order in $\alpha_{s}$. It is the starting point of the evolution from the low scale $1/b_{T}\sim q_{T}$ to the high scale $Q$, which resums the logarithms of $Qb_{T}\sim Q/q_{T}$.
The analogous beam boundary condition was already defined in Eq. (3.6),
$\tilde{b}_{i}[x,\alpha_{s}(b_{0}/b_{T})]\equiv\tilde{B}_{i}(x,b_{T},\mu=b_{0}/b_{T},\nu/Q=1)=\sum_{n=0}\tilde{b}_{i,n}(x)\,\biggl[\frac{\alpha_{s}(b_{0}/b_{T})}{4\pi}\biggr]^{n}\,$, (A.2)
where as shown in Eq. (3.7) the beam boundary coefficients $\tilde{b}_{i,n}(x)$ are computed in terms of PDFs evaluated at $\mu_{f}=b_{0}/b_{T}$.
Note that the beam and soft boundary conditions in Eqs. (A.1) and (A.2) genuinely appear as a low-scale perturbative series in $\alpha_{s}(b_{0}/b_{T})$. The appearance of a perturbative series at the low scale $\mu\sim b_{0}/b_{T}\sim q_{T}$ is a fundamental and universal feature of $q_{T}$ resummation and not specific to a particular resummation framework. (It is also the reason why the beam and soft functions become nonperturbative at sufficiently small $q_{T}\sim 1/b_{T}$, see section 3.3.)
By contrast, the canonical scale of the hard function is $\mu_{H}^{\rm can}=Q$, so its boundary condition is defined as
$\tilde{h}[\alpha_{s}(Q)]\equiv H(Q^{2},\mu=Q)=\sum_{n=0}\tilde{h}_{n}\,\biggl[\frac{\alpha_{s}(Q)}{4\pi}\biggr]^{n}\,$, (A.3)
and thus it genuinely appears as a high-scale perturbative series in $\alpha_{s}(Q)$.
We can now state the difference between NnLL and NnLL more explicitly: At NnLL, the hard, beam, and soft boundary series in Eqs. (A.1), (A.2), and (A.3) are included in Eq. (3.15) up to fixed $\mathcal{O}(\alpha_{s}^{n-1})$ with their respective canonical $\alpha_{s}$. At NnLL, they are instead included up to $\mathcal{O}(\alpha_{s}^{n})$ exactly as they would be at Nn+1LL. Thus, the only difference between NnLL and the full Nn+1LL is that the latter further includes the anomalous dimensions in the evolution factor in Eq. (3.15) to one order higher, see table 3.
@@
Toward higher orders in $q_{T}$ resummation, the contributions from boundary terms become the dominant effect while the anomalous dimensions become less important. The NnLL result is therefore numerically close to the full Nn+1LL result. For example, the SCETlib results at N3LL and N4LL are almost indistinguishable in both central value and scale variation uncertainties [32]. [^12] This applies when using common PDFs. Switching from NNLO to approximate N3LO PDFs, as also formally required for the PDF evolution when switching from N3LL to N4LL, does make a noticeable difference. Furthermore, as the beam and soft boundary coefficients come with a low-scale $\alpha_{s}$, they are much more important than the hard function. We have seen this explicitly in section 4.2.
In DYturbo, following Refs. [109, 110], the low-scale boundary series in the CSS formalism [75], which corresponds to our soft and beam boundary series, is rewritten to superficially appear as high-scale perturbative series as follows (taking the soft function as an example):
$\tilde{s}[\alpha_{s}(b_{0}/b_{T})]$ $=\tilde{s}[\alpha_{s}(Q)]\times\frac{\tilde{s}[\alpha_{s}(b_{0}/b_{T})]}{\tilde{s}[\alpha_{s}(Q)]}=\tilde{s}[\alpha_{s}(Q)]\times\exp\biggl\{\ln\frac{\tilde{s}[\alpha_{s}(b_{0}/b_{T})]}{\tilde{s}[\alpha_{s}(Q)]}\biggr\}$
$=\tilde{s}[\alpha_{s}(Q)]\times\exp\biggl\{-\int_{\alpha_{s}(b_{0}/b_{T})}^{\alpha_{s}(Q)}\!\mathrm{d}\alpha_{s}\,\frac{\mathrm{d}\ln\tilde{s}(\alpha_{s})}{\mathrm{d}\alpha_{s}}\biggr\}$
$=\tilde{s}[\alpha_{s}(Q)]\times\exp\biggl\{-\int_{b_{0}/b_{T}}^{Q}\!\mathrm{d}\ln\mu\,\frac{\mathrm{d}\alpha_{s}}{\mathrm{d}\ln\mu}\,\frac{\mathrm{d}\ln\tilde{s}(\alpha_{s})}{\mathrm{d}\alpha_{s}}\biggr\}\,.$ (A.4)
The first line is clearly trivial, and in the remaining lines the exponential is further cast into a form that resembles the exponential evolution factor in Eq. (3.15). As long as $\tilde{s}(\alpha_{s})$ is consistently included to the same order in $\alpha_{s}$ everywhere it appears, Eq. (A.2) is a strict identity, i.e., there is no formal or numerical difference between the left-hand side and the various versions on the right-hand side. [^13] For the last line in Eq. (A.2), this also requires the running of $\alpha_{s}(\mu)$ to be treated in the same way on both sides of the equation. Otherwise there will be formally higher-order differences which are however irrelevant for our discussion. This is the case at NnLL and NnLL.
However, in the order-counting adopted at NnLL, the last line of Eq. (A.2) is used with the factor $\tilde{s}[\alpha_{s}(Q)]$ considered as a fundamental high-scale boundary series, and akin to the treatment of the hard boundary series at NnLL, it is included to $n$-loop order (i.e. $\tilde{s}_{n}$). At the same time, the exponential factor is considered to be part of the overall evolution factor, which is included to $n$-loop order (akin to the evolution order used at NnLL). As a result, the $\tilde{s}(\alpha_{s})$ boundary series in the exponent is only included to (n1)(n-1)-loop order (i.e. to $\tilde{s}_{n-1}$). This is because the running of $\alpha_{s}$ starts at one loop, so the (n1)(n-1)-loop boundary coefficient $\tilde{s}_{n-1}$ only appears as an $n$-loop contribution as $\beta_{0}\,\tilde{s}_{n-1}$.
However, rewriting the boundary series as in Eq. (A.2) does not change the fundamental fact that it is a low-scale series. This fact is merely hidden on the right-hand side. Reversing the steps in Eq. (A.2) with the NnLL counting applied to the last line on the right-hand side, it results in effectively using the following $n$-loop boundary series on the left-hand side,
$\tilde{s}^{(n)}[\alpha_{s}(Q)]\times\frac{\tilde{s}^{(n-1)}[\alpha_{s}(b_{0}/b_{T})]}{\tilde{s}^{(n-1)}[\alpha_{s}(Q)]}=\sum_{k=0}^{n-1}\tilde{s}_{k}\,\biggl[\frac{\alpha_{s}(b_{0}/b_{T})}{4\pi}\biggr]^{k}+\tilde{s}_{n}\biggl[\frac{\alpha_{s}(Q)}{4\pi}\biggr]^{n}\,$, (A.5)
where the superscripts denote the order in $\alpha_{s}$ to which each series is kept. Hence, due to the mismatch in the treatment of $\tilde{s}(\alpha_{s})$ between prefactor and exponent, the $n$-loop coefficient $\tilde{s}_{n}$ remains multiplied by the high-scale $\alpha_{s}^{n}(Q)$, as indicated in table 3, while its correction to the low scale is postponed to Nn+1LL. By contrast, it is directly included at its correct low scale at both NnLL and full Nn+1LL.
The analogous discussion applies to the beam boundary series, which is rewritten in the same fashion to appear as a high-scale series, including also PDFs evaluated at the high scale, multiplied by a corresponding correction factor that also includes the evolution of the PDFs back down to their natural low scale. At NnLL, the beam boundary terms are included at their correct low scale only to $\tilde{b}_{i,n-1}$, while $\tilde{b}_{i,n}$ is included in terms of $\alpha_{s}$ and PDFs at the high scale.
As a side remark, we comment on the counting of towers of logarithms $\alpha_{s}^{n}L^{m}$ in the perturbative series, with some choice of the large logarithm $L$, which traditionally is often used to define resummation orders. When counting logarithms in the cross section (formally counting $\alpha_{s}L^{2}\sim 1$), NnLL and NnLL capture the same complete logarithmic towers and one more than NnLL. When counting logarithms in the exponent (formally counting $\alpha_{s}L\sim 1$), all three orders capture the same complete logarithmic towers. This does not mean that the orders are equivalent for all purposes. The three orders differ by including different subsets of higher logarithmic towers, which are well defined and numerically relevant. This simply means that their relevant distinction is not (or not easily) visible from a traditional counting of logarithms. Instead, here we define the resummation order (or formal accuracy) at the level of the RGE and its ingredients, which allows us to make a formal distinction between the orders as shown in table 3.
This rather peculiar treatment of the low-scale boundary conditions at NnLL has several consequences. Since $\alpha_{s}(b_{0}/b_{T})/\alpha_{s}(Q)$ can become $\gtrsim 2$ at small $1/b_{T}\sim q_{T}$, and this mismatch is raised to the $n$th power, the above-mentioned dominant effect of the $n$-loop soft and beam boundary coefficients is not yet present at NnLL. For example, the effect of $\tilde{s}_{4}$ and $\tilde{b}_{i,4}$ can be easily suppressed by an order of magnitude or more at N4LL, which is indeed what we find below. Their full impact then only appears at the next order in the disguise of a large evolution effect. Moreover, since they are only included as a high-scale boundary condition like the hard function, also their effect on the shape of the spectrum, which is particularly relevant for extracting $\alpha_{s}$, is not yet present. For these reasons, the N4LL result does not provide a tangible improvement, formally or practically, compared to N4LL. This is confirmed by the comparable scale variations we observed at the spectrum level in the previous subsection. In particular, and regardless of being approximate or exact, it does not provide a reliable approximation of N4LL.
The analysis in Ref. [33] is performed at the so-called approximate N4LL (“N4LLa”) order of Ref. [30], meaning at approximate N4LL. There are two types of approximations involved: First, the 4-loop evolution is formally approximate, because the required 4-loop splitting functions and 5-loop cusp anomalous dimensions are only known approximately. These approximations are also present at N4LL. They are expected to cause more minor effects and are not discussed here further. [^14] Although the available approximations for 4-loop splitting functions are increasingly accurate for phenomenology [111], their effect could in principle be of some relevance and deserves further study. The second type of approximation comes from the fact that the 4-loop beam and soft coefficients, $\tilde{s}_{4}$ and $\tilde{b}_{i,4}$, are completely unknown.
In Refs. [30, 33] the additional uncertainty due to the 5-loop cusp anomalous dimension and the unknown 4-loop beam and soft coefficients are estimated to be $0.1-0.2\%$ for the $q_{T}$ spectrum with a resulting negligible $\pm 0.04\times 10^{-3}$ effect on $\alpha_{s}(m_{Z})$. Since at N4LL the latter are effectively treated like a hard-function contribution, this is consistent with the rather negligible impact of the hard function and cusp anomalous dimension we observe. However, as explained above, this does not provide a realistic estimate of the actual possible impact of the missing 4-loop beam and soft coefficients.
@@
The dominant effect of the beam and soft functions is clearly evident in our TNP studies, where as expected they cause by far the largest uncertainties. From our TNP scanning at N3+1LL, we know that the 3-loop beam and soft functions each cause an effect of order $\sim 1\times 10^{-3}$ on $\alpha_{s}(m_{Z})$. We might expect the corresponding uncertainties due to the 4-loop beam and soft functions to be between two and a few times smaller than that, but certainly not 30 to 40 times smaller. We can obtain a more realistic estimate of the impact of the missing 4-loop beam and soft functions from their associated TNPs at N4+0LL (see appendix B.2). [^15] Ideally, we would use N4+1LL, which we have not yet available, but N4+0LL provides a sufficient approximation for our purposes here. From the analogous TNP scanning at this order we find a combined perturbative uncertainty on $\alpha_{s}(m_{Z})$ due to the missing 4-loop beam and soft functions of $0.75\times 10^{-3}$, as given in table 2, so indeed a reduction by roughly a factor of 2 compared to N3+1LL in line with expectation. Their full effect would appear at N5LL and already by itself can be twice as large as the quoted scale-variation uncertainty at N4LL, providing further evidence that the latter is strongly underestimated.
@@ A.3 Effects due to finite bottom and charm quark masses
Ref. [33] uses a massless 5-flavor description by default. The uncertainties due to neglecting quark-mass effects are estimated by changing to a variable-flavor massless description in various parts of the calculation and by varying the associated threshold scales. From the above discussion, this can only serve as a rough indication of the expected order of magnitude of mass effects but it may or may not reflect the actual size of the effects. Another estimate comes from including the effect of final-state gluon splitting into massive quark pairs. Unfortunately, Ref. [33] provides no further information than that, so it is unclear to us to what extent this provides a complete description of secondary mass effects. The final uncertainty on $\alpha_{s}(m_{Z})$ due to mass effects is taken as the envelope of all observed differences yielding $^{+0.40}_{-0.29}\times 10^{-3}$.
@@ A.4 Nonperturbative model
This model originates from Ref. [94], and while it has traditionally been used in the past, it has several limitations that are potentially important at the high level of precision required here. First, the model does not account for any flavor or $x$ dependence in the nonperturbative TMD parameters, which in general cannot be neglected [99, 31, 100]. As discussed in section 3.3.2, since the fit to real data includes multiple rapidity bins, the nonperturbative TMD model should at minimum allow for an effective rapidity dependence.
@@ B.1 Nonsingular power corrections
We perform a set of Asimov tests following the procedure in section 4.2.3. For simplicity, we do not fit the nonperturbative parameters here, since doing so does not change our conclusions. The nominal perturbative uncertainty for this case obtained in section 4.2.3 is $\Delta_{\rm pert}=_{-0.50}^{+0.47}\times 10^{-3}$. The Asimov data are now defined as the N4LL++NLO1 prediction at $\alpha_{s}(m_{Z})=0.118$, which we fit with three different theory models that differ in the nonsingular terms they include. The results are shown in table 4. In the case where no nonsingular term is included (top row), we find an uncertainty $\Delta_{\rm pert}=0.41\times 10^{-3}$, slightly smaller than the nominal one. Including the LO1 and NLO1 terms (middle and bottom rows), the uncertainty increases again marginally and closely reproduces the nominal one. We conclude that under the assumptions and setup of our pseudodata analysis, the nonsingular contributions do not affect the uncertainty of $\alpha_{s}(m_{Z})$. They can therefore be safely neglected for our purposes. However, as also seen in table 4, neglecting the nonsingular corrections in the theory model induces a substantial bias in the central value of $\alpha_{s}(m_{Z})$, which will necessitate their inclusion in the fit to real data.